Circular No. 8461

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

NOVA IN M31 A rather bright apparent nova in M31 has been discovered indepen-

dently on an R-band CCD frame taken by K. Hornoch (Lelekovice, Czech Republic, 0.35-m reflector) on Jan. 7.891 UT and on unfiltered CCD frames taken by R. Arbour (South Wonston, England, 0.3-m reflector) on Jan. 10.925. Hornoch’s position for the new object: α = 0^{h}42^{m}28^{s}.38, δ = +41^{o}16′36′′.2 (equinox 2000.0), which is 180′′ west and 28′′ north of the center of M31; Arbour provides position end figures 28^{s}.34, 37′′.6 (estimating the object’s magnitude as ∼ 15). Available R-band magnitudes for the nova provided by Hornoch: 2004 Dec. 18.068, [21 (3.5-m WIYN telescope image taken by P. Garnavich); 2005 Jan. 2.821, [19.8; 6.766, 19.2 (prediscovery image taken at Ondřejov by P. Kušnirák); 6.793, 19.4; 7.891, 17.9; 9.724, 15.6; 9.937, 15.27; 10.749, 15.48 (Kušnirák); 10.766, 15.32; 10.786, 15.27; 10.806, 15.30; 10.908, 15.34; 10.926, 15.31. A V -band image by Kušnirák taken on Jan. 10.753 yields mag 15.45.

SUPERNOVAE 2004gq AND 2005A M. Modjaz, R. Kirshner, and P. Challis, Harvard-Smithsonian Center

for Astrophysics; and T. Matheson, National Optical Astronomy Observatory, report that a spectrum (range 350–740 nm) of SN 2005A (cf. IAUC 8459), obtained by E. Falco on Jan. 7.24 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST), shows it to be a type-Ia supernova, one or two weeks before maximum. The supernova expansion velocity, derived from the minimum of Si II (rest 635.5 nm) and adopting the NED recession velocity of 5738 km/s for the host galaxy, is ∼ 14400 km/s. Zerovelocity interstellar Na I D absorption with an equivalent width of ∼ 0.5 nm is detected in a spectrum of SN 2005A taken on Jan. 8.24 by P. Berlind, indicating gas along the line-of-sight in our Galaxy and thus suggesting reddening by dust. Schlegel et al. (1998, Ap.J. 500, 525) estimate a Galactic reddening of E(B−V) = 0.03 mag along the line-of-sight to the host galaxy (NGC 958).

Modjaz et al. add that a spectrogram of SN 2004gq (cf. IAUC 8452), obtained by Falco on Jan. 7.29 UT, shows it to be a type-Ib supernova. The spectrum exhibits conspicuous lines of He I (rest 447.1, 587.6, 667.8, 706.5 nm) and is very similar to spectra of SN 1984L (Harkness et al. 1987, Ap.J. 317, 355), ∼ 3 weeks after maximum. SN 2004gq was previously classified (IAUC 8404) as a type-Ic supernova.

2005 January 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green